OpenStudy (anonymous):

. How many zeros does this function have? y = 5x2 + 10x - 3 2 1 0

OpenStudy (anonymous):

check that \(b^2-4ac=10^2-4\times 5\times (-3)\) is greater than zero (positive) that tells you it has two real zeros